OpenStudy (anonymous):

when a ball is dropped off a 50 foot building, it is h feet above the ground after t seconds where t= sqrt(50-h)/4 how long for the ball to hit ground how far above the ground will the ball be after 1 second

OpenStudy (anonymous):

t= sqrt(50-h)/4 When it hits the ground, h=0 so: t= sqrt(50)/4 = 1.767767 seconds. after one second, t=1, so 1 = sqrt (50-h) /4 sqrt (50-h) = 4 50 -h = 16 h=34 feet.

OpenStudy (anonymous):

you are amazing

OpenStudy (anonymous):

I only have 8 fans...

OpenStudy (anonymous):

i became a fan but it didnt do anything

OpenStudy (anonymous):

my 1st time on this site

OpenStudy (anonymous):

Doesn't matter, I don't come here too often either.